OpenStudy (anonymous):

The population in millions of a certain country can be approximated by the function: p(x) = 50(1.03)^x Where x is the number of years since 2000. in which year will the population reach 100 million ?

OpenStudy (anonymous):



OpenStudy (anonymous):



OpenStudy (anonymous):

The answer choices are: 2023, 2066. 2132, and 2002.

OpenStudy (anonymous):

What did I do wrong ??

OpenStudy (faiqraees):

100 =50(1.03)^x 2=1.03^x lg2 =x lg1.03 x = lg2/lg1.03 x=.....?

OpenStudy (anonymous):

x = 23.45 ..... which would be 2023 ...@FaiqRaees

OpenStudy (anonymous):

Thank you @FaiqRaees

OpenStudy (faiqraees):

np